OpenStudy (anonymous):

A Country's population in 1990 was 123 million. In 2002 it was 128 million. Estimate the population in 2013 using the exponential growth formula. Round your answer to the nearest million

OpenStudy (anonymous):

I got 132.762, so rounded to the nearest million that would be 133

OpenStudy (anonymous):

First find k : p=A^e (kt) 128=123e^(kt) 1.04065=e^(12k) In(1.04065)=12k k=In(1.04065)/12=0.0033 To solve; p=128e^(0.0033)(1)=128(1.0033) =128.423~128 million

OpenStudy (anonymous):

P2 = P1 exp(kt) exponential growth 123 = 128 exp(12 k) 12 k = ln(123/128) find k P3 = 128 exp(11 k)